# Introspectively Envy-Free and Efficient Allocation of Indivisible Mixed Manna

### v2

@@ Appendix A $\mathrm{EF}1$ and $\mathrm{PO}$ for Indivisible Goods
There exists a point $\overline{w}^{*}\in\Delta_{n-1}$ such that for each $i\in[n]$, there is an allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\eta))$ such that $i\in\operatorname*{arg\,min}_{j\in[n]}p(A^{(i)}_{j})$ where $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\eta))$.
The subsequent argument to prove the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocation also closely follows Section 6. Let $\overline{w}^{*}\in\Delta_{n-1}$ as defined in Theorem 25. Define $p^{*}=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\eta))$. Write $\overline{L}_{i}\coloneqq\left\{s\in[\widehat{m}]\mid\widehat{v}_{i}(s)/(\overline{w}^{*}_{i}+\eta)>\widehat{v}_{j}(s)/(\overline{w}^{*}_{j}+\eta)\text{ for all }j\neq i\right\}$, for each $i\in[n]$, to denote the set of items that must be allocated to agent $i$ in any allocation in $\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\eta))$. Denote by $\overline{\mathcal{L}}=(\overline{L}_{1},\ldots,\overline{L}_{n})$ the partial allocation formed by these bundles. Denote by $\overline{U}\coloneqq[\widehat{m}]\setminus\left(\cup_{j=1}^{n}\overline{L}_{j}\right)$.
Define a bipartite graph $\overline{H}=([n]\cup\overline{U},\overline{E})$ of $\overline{G}$ with $[n]$ and $\overline{U}$ being the two partitions such that $(i,j)\in\overline{E}\subseteq[n]\times\overline{U}$ if and only if $p^{*}_{j}=\widehat{v}_{i}(j)/(\overline{w}^{*}_{i}+\eta)$. The vertices adjacent to any vertex $i$ in $\overline{H}$ will be denoted by $\overline{\Gamma}(i)$. As in Section 6, the subsequent results hold.
@@
Define $\overline{O}^{*}\coloneqq\{(B_{1},B_{2},\ldots,B_{n})\in\Pi_{n}([\widehat{m}])\mid\text{ for all }i\in[n],\overline{L}_{i}\subseteq B_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)\}$. Similar to Lemma 12, the following lemma holds.
@@
Any allocation $\mathcal{A}\in\overline{O}^{*}$ if and only if $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\eta))$.
@@
We now define some key notations. These notations are analogous to Section 6.2, yet crucially differ from it. We begin by defining the max-min price threshold,
$\overline{\tau}\coloneqq\max_{(B_{1},B_{2},\ldots,B_{n})\in\overline{O}^{*}}\min_{j\in[n]}p^{*}(B_{j}),$
unlike $\tau$ (Equation 19), $\overline{\tau}$ has a max-min formulation. Define set $\overline{\mathcal{M}}^{*}\coloneqq\{(B_{1},\ldots,B_{n})\in\overline{O}^{*}\mid\min_{j\in[n]}p^{*}(B_{j})=\overline{\tau}\}$ as the collection of all max-min allocations. Additionally, for all agents $i\in[n]$ and subsets $S$, we define
$p^{-}_{i}(S)=\min\limits_{T\subseteq\overline{\Gamma}(i),|T|\leq 1}p^{*}(S\triangle T).$
We now define the following potential for any allocation $\mathcal{B}=(B_{1},B_{2},\ldots,B_{n})$,
$\varphi^{-}(\mathcal{B})\coloneqq|\{i\in[n]\mid p_{i}^{-}(B_{i})\leq\overline{\tau}\}|.$
The subsequent goal will be to prove the following theorem.
@@
Assuming Theorem 27, we will now establish the existence of $\mathrm{XEF}1$ and $\mathrm{PO}$ allocations.
@@
Every mixed manna instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, with additive valuations, admits an allocation that is $\mathrm{XEF}1$ and $\mathrm{PO}$.
@@
Consider the weight vector $\overline{w}^{*}\in\Delta_{n-1}$ whose existence is implied by Theorem 25 and let $p^{*}=\allowbreak\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\eta))$. Let $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ be the allocation whose existence is implied by Theorem 27.
Note that $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}\subseteq\overline{O}^{*}$, therefore by Proposition 4, we get $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\eta))$. Now we will show that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ for the instance $\overline{\mathcal{I}}$. By Theorem 27, we have $p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}$ for all $j\in[n]$. Since, $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ we also have $\overline{\tau}=\min_{j=1}^{n}p^{*}(A^{*}_{j})$. Hence, for any pair of agents $i,j\in[n]$, we have
$\min_{T\subseteq\overline{\Gamma}(j),|T|\leq 1}p^{*}(A^{*}_{j}\triangle T)=p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}\leq p^{*}(A^{*}_{i}).$ (34)
Note that, for each $t\in A^{*}_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)$ we have, $p^{*}_{t}=\widehat{v}_{i}(t)/(w_{i}+\eta)$. In addition to this, for each $k\in A^{*}_{j}\Delta T$, we have $p^{*}_{k}=\max_{a=1}^{n}\widehat{v}_{a}(k)/(w_{a}+\eta)\geq\widehat{v}_{i}(k)/(w_{i}+\eta)$. Substituting this in Equation 34, we get
$\frac{1}{w_{i}+\eta}\cdot\min_{T\subseteq\overline{\Gamma}(j),|T|\leq 1}\widehat{v}_{i}(A^{*}_{j}\triangle T)\leq\min_{T\subseteq\overline{\Gamma}(j),|T|\leq 1}p^{*}(A^{*}_{j}\triangle T)\leq p^{*}(A^{*}_{i})=\frac{1}{w_{i}+\eta}\cdot\widehat{v}_{i}(A^{*}_{i}).$ (35)
Using the fact that $w_{i}+\eta>0$, we get that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ for $\widehat{\mathcal{I}}$, i.e., for each agent $i,j\in[n]$, there is a subset $T\subseteq\overline{\Gamma}(j)\subseteq[\widehat{m}$] with $|T|\leq 1$ such that $\widehat{v}_{i}(A^{*}_{i})\geq\widehat{v}_{i}(A^{*}_{j}\triangle T)$.
Hence, the allocation $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ for $\overline{\mathcal{I}}$ and $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\eta))$. Finally, Lemmas 21 and 22 implies that the allocation $(A^{*}_{1}\setminus\{m+1\},A^{*}_{2}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{PO}$ and $\mathrm{XEF}1$ for the given instance $\mathcal{I}$, thereby completing the proof. ∎
@@ A.1 Proof of Theorem 27
Input : Allocation $\mathcal{S}=(S_{1},\dots,S_{n})\in\overline{\mathcal{M}}^{*}$ (i.e., $\min_{j}p^{*}(S_{j})=\overline{\tau}$) with $\varphi^{-}(\mathcal{S})<n$.
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\overline{\mathcal{M}}^{*}$ satisfying $\varphi^{-}(\widehat{\mathcal{S}})>\varphi^{-}(\mathcal{S})$
@@
13  1exSet $X\subseteq\mathrm{ch}(i)$ to be an item subset that satisfies: (i) $p^{-}_{i}(\widehat{S}_{i}\triangle X)\leq\overline{\tau}$, (ii) $p^{*}(\widehat{S}_{i}\triangle X)\geq\overline{\tau}$, and (iii) $p^{*}(\widehat{S}_{i}\triangle\{t\})<p^{*}(\widehat{S}_{i})$ for each $t\in X$. {Existence of $X$ is guaranteed via Lemma 29.}
14
@@
23
@@
The proofs of the following four lemmas Lemma 29-31 exactly mimic the proofs of Lemma 14-16 and hence are omitted.
@@
The corollary below follows from Lemma 15 and Lemma 16.
@@
We finally prove Theorem 27.

### v3

@@ Appendix A $\mathrm{EF}1$ and $\mathrm{PO}$ for Indivisible Goods
There exists a point $\overline{w}^{*}\in\Delta_{n-1}$ with the property that, for each $i\in[n]$, we have an allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$ wherein $i\in\operatorname*{arg\,min}_{j\in[n]}p^{*}(A^{(i)}_{j})$; here $p^{*}=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$.
The subsequent argument to prove the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocation also closely follows Section 6. Let $\overline{w}^{*}\in\Delta_{n-1}$ be the weight vector whose existence is guaranteed in Theorem 25, and let $p^{*}=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$ be the associated optimal prices.
For each agent $i\in[n]$, write $\overline{L}_{i}$ to denote the set of items that must be allocated to agent $i$ in any allocation in $\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$, i.e., $\overline{L}_{i}\coloneqq\left\{s\in[\widehat{m}]\mid\frac{\widehat{v}_{i}(s)}{(\overline{w}^{*}_{i}+\overline{\eta})}>\frac{\widehat{v}_{j}(s)}{(\overline{w}^{*}_{j}+\overline{\eta})}\text{ for all }j\neq i\right\}$. Denote by $\overline{\mathcal{L}}=(\overline{L}_{1},\ldots,\overline{L}_{n})$ the partial allocation formed by these bundles and let $\overline{U}\coloneqq[\widehat{m}]\setminus\left(\cup_{j=1}^{n}\overline{L}_{j}\right)$.
Next, we consider a bipartite graph $\overline{H}=([n]\cup\overline{U},\overline{E})$, with $[n]$ and $\overline{U}$ as the left and right part, respectively. Edge $(i,t)\in\overline{E}\subseteq[n]\times\overline{U}$ is included in the bipartite graph if and only if $p^{*}_{t}=\frac{\widehat{v}_{i}(t)}{\overline{w}^{*}_{i}+\overline{\eta}}$. The vertices adjacent to any vertex $x$ in $\overline{H}$ will be denoted by $\overline{\Gamma}(x)$. Arguments similar to the ones used in Section 6 establish the following lemma for $\overline{H}$.
@@
Define collection of complete allocations $\overline{O}^{*}\coloneqq\left\{(B_{1},\ldots,B_{n})\in\Pi_{n}([\widehat{m}])\mid\overline{L}_{i}\subseteq B_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)\text{ for each }i\in[n]\right\}$. Analogous to Lemma 12, we have the following result for $\overline{O}^{*}$.
@@
Any allocation $\mathcal{B}\in\overline{O}^{*}$ if and only if $\mathcal{B}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$.
@@
We now define some key constructs for the rest of the analysis. Note that while the high-level structure of the proof here is similar to Section 6.2, multiple constructs differ from the previous ones. In particular, the price threshold $\overline{\tau}$ (defined below) has a max-min formulation in contrast to $\tau$ defined Equation 19.
$\overline{\tau}\coloneqq\max_{(B_{1},\ldots,B_{n})\in\overline{O}^{*}}\ \ \min_{i\in[n]}p^{*}(B_{i}).$
Define set $\overline{\mathcal{M}}^{*}\coloneqq\{(B_{1},\ldots,B_{n})\in\overline{O}^{*}\mid\min_{j\in[n]}p^{*}(B_{j})=\overline{\tau}\}$ as the collection of all max-min allocations. Additionally, for each agent $i\in[n]$ and subsets $S$, we define
$p^{-}_{i}(S)\coloneqq\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(i),\\ |T|\leq 1\end{subarray}}p^{*}(S\triangle T).$
We also define, for each allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$, the potential $\varphi^{-}(\mathcal{B})$, which counts the number of agents whose bundle prices can be toggled to $\overline{\tau}$ or below.
$\varphi^{-}(\mathcal{B})\coloneqq\left|\{i\in[n]\mid p_{i}^{-}(B_{i})\leq\overline{\tau}\}\right|.$
The following theorem is a key result of this section, and it shows that there exists a max-min allocation $\mathcal{A}^{*}$ wherein the bundle prices of all the agents can be toggled to $\overline{\tau}$ or below.
@@
Before proving Theorem 27 (in Section A.1 below), we use this result to establish the main result of this section, the existence of $\mathrm{XEF}1$ and $\mathrm{PO}$ allocations.
@@
Every fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, with indivisible mixed manna and additive valuations, admits an allocation that is both extrospectively envy-free up to one item ($\mathrm{XEF}1$) and Pareto efficient ($\mathrm{PO}$).
@@
Consider the weight vector $\overline{w}^{*}\in\Delta_{n-1}$ whose existence is implied by Theorem 25 and let $p^{*}=\allowbreak\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$. Let $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ be the allocation whose existence is implied by Theorem 27.
First, note that $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}\subseteq\overline{O}^{*}$, therefore by Proposition 4, we get $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$.
Next, we will show that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ for the instance $\overline{\mathcal{I}}$. By Theorem 27, we have $p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}$ for all $j\in[n]$. Since, $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ we also have $\overline{\tau}=\min_{i\in[n]}\ p^{*}(A^{*}_{i})$. Hence, for any pair of agents $i,j\in[n]$, we have
$\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}p^{*}(A^{*}_{j}\triangle T)=p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}\leq p^{*}(A^{*}_{i}).$ (40)
Recall that $\mathcal{A}^{*}\in\overline{O}^{*}$ and, hence, $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$ (Proposition 4). Therefore, for each item $t\in A^{*}_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)$ it holds that $p^{*}_{t}=\widehat{v}_{i}(t)/(\overline{w}^{*}_{i}+\overline{\eta})$. Furthermore, for each item $s\in A^{*}_{j}\Delta T$, we have $p^{*}_{s}=\max_{a\in[n]}\widehat{v}_{a}(s)/(\overline{w}^{*}_{a}+\overline{\eta})\geq\widehat{v}_{i}(s)/(\overline{w}^{*}_{i}+\overline{\eta})$. Using these bounds with Equation 40, we obtain
$\frac{1}{\overline{w}^{*}_{i}+\overline{\eta}}\ \min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}\widehat{v}_{i}(A^{*}_{j}\triangle T)\leq\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}p^{*}(A^{*}_{j}\triangle T)\leq p^{*}(A^{*}_{i})=\frac{1}{\overline{w}^{*}_{i}+\overline{\eta}}\ \widehat{v}_{i}(A^{*}_{i}).$ (41)
Since $(\overline{w}^{*}_{i}+\overline{\eta})>0$, we get that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ in the instance $\widehat{\mathcal{I}}$: For each pair of agents $i,j\in[n]$, there exists a a subset $T\subseteq\overline{\Gamma}(j)\subseteq[\widehat{m}$] of size at most one ($|T|\leq 1$) such that $\widehat{v}_{i}(A^{*}_{i})\geq\widehat{v}_{i}(A^{*}_{j}\triangle T)$.
Therefore, we have that the allocation $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ in $\overline{\mathcal{I}}$ and $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$. Finally, Lemmas 21 and 22 imply that the allocation $(A^{*}_{1}\setminus\{m+1\},A^{*}_{2}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{XEF}1$ and $\mathrm{PO}$ in the given instance $\mathcal{I}$. This completes the proof of the theorem. ∎
@@ A.1 Proof of Theorem 27
Input : Allocation $\mathcal{S}=(S_{1},\dots,S_{n})\in\overline{\mathcal{M}}^{*}$ (i.e., $\min_{i}p^{*}(S_{i})=\overline{\tau}$) with $\varphi^{-}(\mathcal{S})<n$.
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\overline{\mathcal{M}}^{*}$ satisfying $\varphi^{-}(\widehat{\mathcal{S}})>\varphi^{-}(\mathcal{S})$.
@@
7Write $R$ to denote the tree in bipartite graph (forest) $\overline{H}$ rooted at $r$. For each node $x$ in the tree $R$, let $\mathrm{par}(x)$ denote the parent of $x$. In addition, $\mathrm{ch}(x)=\Gamma(x)\setminus\{\mathrm{par}(x)\}$ are the children of $x$ in the tree. Also, for each agent $i$ in the tree, $\overline{\mathrm{ch}}(i)=\cup_{t\in\mathrm{ch}(i)}\ \mathrm{ch}(t)$ denotes the agents that are immediate descendants of $i$.
@@
13  1ex
14  Set $X\subseteq\mathrm{ch}(i)$ to be an item subset that satisfies: (i) $p^{-}_{i}(\widehat{S}_{i}\triangle X)\leq\overline{\tau}$, (ii) $p^{*}(\widehat{S}_{i}\triangle X)\geq\overline{\tau}$, and (iii) $p^{*}(\widehat{S}_{i}\triangle\{t\})<p^{*}(\widehat{S}_{i})$ for each $t\in X$. {Existence of $X$ is guaranteed via Lemma 29.}
15
@@
17      If item $t\in\widehat{S}_{i}$, then set $a$ to be any agent in $\mathrm{ch}(t)\neq\emptyset$. Otherwise, if $t\notin\widehat{S}_{i}$, then let $a\in\mathrm{ch}(t)$ be the agent such that $t\in\widehat{S}_{a}$. {Such an agent $a\in\overline{\mathrm{ch}}(i)$ always exists.}
@@
24
25return Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})$.
@@
The proofs of the following three lemmas (Lemmas 29 to 31) follow along the lines of the proofs of Lemmas 14 to 16, and hence are omitted.
@@
The corollary below follows from Lemmas 30 and 31.
@@
The $\mathrm{XEF}1$AugmentTree Algorithm (Algorithm 2) always terminates and returns a complete allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})\in\mathcal{M}^{*}$.
We now establish Theorem 27.
@@
Appendix B Scope for $\mathrm{EF}1$
This section provides an example to show that, starting from Theorem 10, the best one can hope for is an $\mathrm{IEF}1$—and not an $\mathrm{EF}1$—guarantee.
Consider an instance with three agents ($n=3$) and three items, $\{g_{1},g_{2},g_{3}\}$. The valuations of the items (all goods) are listed in Table 1.
$g_{1}$ $g_{2}$ $g_{3}$
$v_{1}(\cdot)$ 30 1 1
$v_{2}(\cdot)$ 30 1 1
$v_{3}(\cdot)$ 30 3 3
Table 1: Valuations of the three agents for the three goods.
As detailed in Section 4.1, we additively perturb the given valuations $v_{i}(t)$s—by random $\varepsilon_{i,t}$s—to construct the values $\widehat{v}_{i}(t)$s in non-degenerate instance. In particular, let $\varepsilon_{1,1}$, $\varepsilon_{2,1}$, and $\varepsilon_{3,1}$ be the perturbations for the first good, $g_{1}$, among the three agents, respectively. Also, let $\eta>0$ be as defined in Equation 4.
Now, consider an weight vector $w^{*}=(w^{*}_{1},w^{*}_{2},w^{*}_{3})$ whose nonnegative components sum up to one and satisfy
$(w^{*}_{1}+\eta)\left(30-\varepsilon_{1,1}\right)=(w^{*}_{2}+\eta)\left(30-\varepsilon_{2,1}\right)=(w^{*}_{3}+\eta)\left(30-\varepsilon_{3,1}\right)$ (42)
Since $\varepsilon_{i,t}$s and $\eta$ are sufficiently small, there exist such $w^{*}_{1},w^{*}_{2},w^{*}_{3}\approx 1/3$.
[…]
$g_{2}$ $g_{3}$
$p^{*}$ $\left(w^{*}_{3}+\eta\right)\left(3-\varepsilon_{3,2}\right)$ $\left(w^{*}_{3}+\eta\right)\left(3-\varepsilon_{3,3}\right)$
We next note that the above-mentioned weight vector $w^{*}\in\Delta_{n-1}$ and optimal prices $p^{*}$ satisfy Theorem 10. Here, in any optimal allocation $(A^{*}_{1},A^{*}_{2},A^{*}_{3})\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$, goods $g_{2}$ and $g_{3}$ are necessarily assigned to agent $3$. That is, for any optimal allocation $(A^{*}_{1},A^{*}_{2},A^{*}_{3})$ it holds that $\{g_{2},g_{3}\}\subseteq A^{*}_{3}$. The good $g_{1}$ can be allocated to any one of the three agents while maintaining optimality. [^16] The auxiliary good introduced in the non-degenerate instance has a sufficiently low value and, hence, its allocation does not impact the arguments here.
These observations ensure that $w^{*}$ and $p^{*}$ conform to the guarantee in Theorem 10. In particular, each agent $i\in[3]$ can be made price envy-free by assigning $g_{1}$ to it. Specifically, for each $i\in[3]$, including $g_{1}$ in $A^{(i)}_{i}$ gives us an optimal allocation $\mathcal{A}^{(i)}$ wherein $i\in\operatorname*{arg\,max}_{k\in[3]}\ p^{*}(A^{(i)}_{k})$. For instance, for $A^{(1)}_{1}=\{g_{1}\}$ along with $A^{(1)}_{2}$$\emptyset$ and $A^{(1)}_{3}=\{g_{2},g_{3}\}$, we have $p^{*}(A^{(1)}_{1})\geq 9>p^{*}(A^{(1)}_{2})$ and $p^{*}(A^{(1)}_{1})\geq 9>p^{*}(A^{(1)}_{3})$.
Also, observe that this allocation $\mathcal{A}^{(1)}=(A^{(1)}_{1},A^{(1)}_{2},A^{(1)}_{3})$ is $\mathrm{IEF}1$ (under the given valuations $v_{i}$s) – agent $1$ is envy-free under $\mathcal{A}^{(1)}$ and agents $2$ and $3$ can (hypothetically) include good $g_{1}$ into their own bundle to eliminate their existing envy.
However, in this setting, no allocation in $\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ is $\mathrm{EF}1$: As mentioned previously, under any optimal allocation $\mathcal{A}=(A_{1},A_{2},A_{3})$ goods $g_{2}$ and $g_{3}$ are necessarily assigned to agent $3$. Now, either the first or the second agent will not receive $g_{1}$ in $\mathcal{A}$. For such an agent $i\in\{1,2\}$, we have $v_{i}(A_{i})=0<v_{i}(A_{3})$ and the envy persists even after removal of any good from $A_{3}$, since $v_{i}(A_{3}\setminus\{g\})\geq 1$ for each $g\in A_{3}$.
Therefore, starting with a weight vector $w^{*}$ and prices $p^{*}$ as in Theorem 10, it is not necessary that we can identify an $\mathrm{EF}1$ allocation in $\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$.